# Operational research approaches and mathematical models for kidney exchange: A literature survey and empirical evaluation

### v1

@@ 3.5 Edge Formulation
In EF-CYCLE, we introduce a binary decision variable $x_{uv}$ for every arc $(u,v)\in\mathcal{A}_{\mathcal{R}}$, taking value $1$ if arc $(u,v)$ is selected, and value 00 otherwise. Moreover, we define $\mathcal{P}_{K-1}$ as the set of all maximal cycle-feasible paths, where a maximal cycle-feasible path $p$ is a simple path in $\mathcal{G}$ of length $K-1$. We denote the $i$th vertex of path $p$ by $p_{i}$ for $i=1,\ldots,K$. Note that it is only possible to select all the arcs along a maximal cycle-feasible path if the arc $(p_{K},p_{1})$ from the head of the path to the tail of the arc (if it exists) is selected as well. This leads to the following definition of EF-CYCLE:
[…]
@@
The objective function (16) maximises the total weight and constraints (17) enforce that all RDPs are involved in at most one cycle. Moreover, constraints (18) are the flow conservation constraints ensuring that whenever a unit of flow enters a vertex, it must leave the vertex too (i.e., if a recipient receives a kidney, their paired donor should donate a kidney as well), and constraints (19) are the long-cycle elimination constraints, which forbid all cycles with length exceeding $K$ [^4] Strictly speaking, the term $x_{p_{K},p_{1}}$ should be multiplied by $\mathbbm{1}_{(p_{K},p_{1})\in\mathcal{A}_{\mathcal{R}}}$ to ensure that it is only subtracted if the underlying arc $(p_{K},p_{1})$ exists. For brevity, we omit these coefficients in this model and in the models that follow, in analogous scenarios..
Example 1.
(continued) Consider the instance presented in Figure 3 for $K=3$. EF-CYCLE contains $6$ variables, one for each arc in $\mathcal{A}_{\mathcal{R}}$. For instance, cycle $\left\langle 2,3,4,2\right\rangle$ is obtained by setting $x_{23}=x_{34}=x_{42}=1$. Constraint (19) for maximal cycle-feasible path $\left\langle 1,2,3\right\rangle$ is required to forbid cycle $\left\langle 1,2,3,4,1\right\rangle$.
In the following, we discuss two chain models based on EF-CYCLE, which we call EF-CHAIN-EXP and EF-CHAIN-MTZ. The former is strongly inspired by models introduced by Constantino et al., (2013) and Anderson et al., (2015), whereas the latter is a variant of a model introduced by Mak-Hau, (2017). Mak-Hau, (2017) combined their version of EF-CHAIN-MTZ with both EF-CYCLE and EEF-CYCLE, leading to the “exponential-sized SPLIT formulation” and the “polynomial-sized SPLIT formulation”. The main challenge in adapting EF-CYCLE to a chain model is that not only all chains of length more than $L$ must be excluded, but also that all cycles must be forbidden (following our definition of chain models). To deal with this, EF-CHAIN-EXP contains an exponential number of constraints. On the other hand, EF-CHAIN-MTZ borrows ideas from the Miller-Tucker-Zemlin model for the travelling salesman problem (Miller et al., , 1960), resulting in a model of polynomial size.
In both EF-CHAIN-EXP and EF-CHAIN-MTZ, we introduce a binary decision variable $y_{uv}$ for every arc $(u,v)\in\mathcal{A}$ (including not only $\mathcal{A}_{\mathcal{R}}$, but also $\mathcal{A}_{\mathcal{N}}$ and $\mathcal{A}_{\tau}$), taking value $1$ if arc $(u,v)$ is selected, and value 00 otherwise. For EF-CHAIN-EXP, we define $\mathcal{P}^{\prime}_{L-1}$ as the set of minimal chain-infeasible paths, where a minimal chain-infeasible path $p$ is a simple path in $\mathcal{G}$ of length $L-1$ restricted to the vertices in $\mathcal{R}$. Note that it is not possible to use all the arcs along a minimal chain-infeasible path, as this would result in a chain of length exceeding $L$ (after initialising the chain by an NDD and terminating the chain in $\tau$). Moreover, we define $\mathcal{C}_{\leq L-1}$ as the set of cycles of length at most $L-1$. EF-CHAIN-EXP is then as follows:
[…]
@@
The objective function (21) maximises the total weight, constraints (22) and (23) enforce that all NDDs and RDPs, respectively, are involved in at most one chain, and constraints (24) ensure flow conservation. Furthermore, constraints (25) forbid all chains of length more than $L$. Note that these constraints automatically also forbid all cycles of length at least $L$. Therefore, constraints (26) are added to forbid all the other cycles (i.e., those with length at most $L-1$). We mention that Constantino et al., (2013) and Anderson et al., (2015) consider a slightly different setting. In particular, the model proposed by Constantino et al., (2013) does not contain cycle-elimination constraints, whereas the model by Anderson et al., (2015) does not eliminate the long chains.
Example 1.
(continued) Consider the instance presented in Figure 3 for $L=4$. EF-CHAIN-EXP contains $14$ variables, one for each arc in $\mathcal{A}$. For instance, chains $\left\langle 6,2,3,4,\tau\right\rangle$ and $\left\langle 5,1,\tau\right\rangle$ are obtained by setting $y_{62}=y_{23}=y_{34}=y_{4\tau}=y_{51}=y_{1\tau}=1$. Several long chains and cycles must be forbidden. For example, chain $\left\langle 5,1,2,3,4,\tau\right\rangle$ is excluded by constraint (25) for minimal chain-infeasible path $\left\langle 1,2,3,4\right\rangle$, and cycle $\left\langle 2,3,4,2\right\rangle$ is excluded by the constraint (26) corresponding to this cycle.
On the other hand, the model EF-CHAIN-MTZ does not involve sets $\mathcal{P}^{\prime}_{L-1}$ and $\mathcal{C}_{\leq L-1}$, but does contain an additional “timestamp” variable $t_{v}$ for every RDP $v\in\mathcal{R}$. Using these variables, EF-CHAIN-MTZ is as follows [^5] The term $(L-3)y_{vu}$ in constraints (29) is omitted when $L\leq 2$.):
[…]
@@
Constraints (29) and (30) together ensure that the maximum chain length is respected and that all cycles are forbidden. Indeed, for each arc $(u,v)\in\mathcal{A}_{\mathcal{R}}$, the associated constraint from (29) imposes that $t_{v}\geq t_{u}+1$ if the arc is selected (i.e., if $y_{uv}=1$), whatever the value of $y_{vu}$. This means that the $t$-variables can really be interpreted as timestamps. Moreover, the term $(L-3)y_{vu}$ (which is only added when the underlying arc $(v,u)$ exists and $L\geq 3$), is optional, but strengthens the linear relaxation. Indeed, for every $(u,v)\in\mathcal{A}_{\mathcal{R}}$, constraints (29) for $(u,v)$ and $(v,u)$ together imply that $y_{uv}+y_{vu}\leq 1$, which can be used to deduce that $y_{vu}=0$ when $y_{uv}=1$, which in turn implies that $t_{v}=t_{u}+1$. Mak-Hau, (2017) presented an equivalent formulation that includes additional timestamp variables $t_{v}$ for all $v\in\mathcal{N}\cup\{\tau\}$.
Example 1.
(continued) Consider again the instance presented in Figure 3 for $L=4$. Compared to EF-CHAIN-EXP, model EF-CHAIN-MTZ contains $4$ additional timestamp variables, one per RDP. In return, fewer constraints are required to forbid long chains and cycles. To obtain chains $\left\langle 6,2,3,4,\tau\right\rangle$ and $\left\langle 5,1,\tau\right\rangle$, we must have $t_{1}=1$, $t_{2}=1$, $t_{3}=2$ and $t_{4}=3$.
@@
Regarding model-related improvements, Mak-Hau, (2017) proposed reducing the model size by removing vertices and arcs that cannot appear in any solution respecting the cardinality constraints. That is, for EF-CYCLE we can remove all vertices/arcs that cannot appear in any cycle of length at most $K$, and in EF-CHAIN we can remove all vertices/arcs that cannot appear in any chain of length at most $L$. For completeness, we present these pre-processing algorithms in Appendix A.5. Furthermore, for EF-CHAIN-EXP and EF-CHAIN-MTZ, it is also possible to omit the variables $y_{v\tau}$ for arcs $(v,\tau)\in\mathcal{A}_{\tau}$, assuming that all weights $w_{v\tau}$ for arcs $(v,\tau)\in\mathcal{A}_{\tau}$ are nonnegative, which we describe in detail in Appendix A.6.
@@
Moreover, note that the number of constraints in EF-CYCLE and EF-CHAIN-EXP is exponential in $K$ and $L$, respectively, which can be problematic. This concerns constraints (19) for EF-CYCLE and constraints (25) and (26) for EF-CHAIN-EXP. Nevertheless, Constantino et al., (2013) and Mak-Hau, (2017) showed that enumerating all constraints is a viable approach in certain cases (i.e., when the compatibility graph is sufficiently small and sparse and $K$ and $L$ are sufficiently small too). However, in the experiments by Constantino et al., (2013), there were for example already too many constraints for EF-CYCLE for high-density instances with $|\mathcal{R}|=20$ and $K\geq 5$. Alternatively, one can apply constraint generation, in which the problematic constraints are initially relaxed, after which one iteratively only adds the constraints that have been violated by previous incumbent solutions until an optimal solution is found. This was for example applied by Abraham et al., (2007), Anderson et al., (2015), Mak-Hau, (2018) and Delorme et al., (2023).
@@
Furthermore, we mention that research has been done on alternatives for some of the problematic constraints mentioned above. In Appendix A.2 we discuss two of the alternatives that were proposed for constraints (19) in EF-CYCLE and one alternative that was proposed for constraints (26) in EF-CHAIN-EXP. However, our preliminary computational experiments indicated that these alternatives either performed similarly to (or worse than) the versions that we presented above. Furthermore, we note that in theory, when EF-CHAIN-EXP is combined with a cycle model, we only need the constraints (26) that forbid cycles of length more than $K$. This would allow us to reduce the number of constraints, at the cost of introducing some symmetry (as some cycles can then be obtained through either the cycle component of the model or the chain component of the model). For simplicity, we do not consider this in our computational experiments.
@@
Last, as mentioned in Section 3.2, it is possible to model cycles and chains concurrently with a single set of variables using ideas similar to those underlying EF-CYCLE and EF-CHAIN-EXP. Such a hybrid model has the potential advantage that its size may be smaller than that of the combined model composed of EF-CYCLE and EF-CHAIN-EXP. However, whereas in a combined model with two separate sets of variables it is relatively easy to forbid all cycles of length more than $K$ without excluding any feasible chain, while also forbidding all chains of length more than $L$ without excluding any feasible cycle, this is a real challenge for a hybrid model with a single set of variables. Indeed, Constantino et al., (2013) present a model that only applies when $L\leq K+1$, as their constraints that eliminate long cycles also forbid feasible chains, otherwise. Alternatively, Anderson et al., (2015) avoided part of the problem by presenting a model for the case where $L=\infty$. They also explained how their model could be extended to the case where $L$ is finite. However, their underlying idea is more similar to the idea behind the chain model EEF-CHAIN-EXP, which we present in Section 3.6. On the other hand, our new model, denoted by EF-HYBRID, fits more closely the paradigm of the models EF-CYCLE and EF-CHAIN-EXP that were presented earlier in this section. However, our experiments showed that EF-HYBRID does not perform well (see Section 4.2.1). Therefore, we present this model in Appendix A.3.
@@ 3.6 Extended Edge Formulation
The main downsides of the models presented in Sections 3.3-3.5 are the exponential number of variables or constraints. On the other hand, the model EEF-CYCLE, introduced by Constantino et al., (2013), can be seen as an extended formulation based on EF-CYCLE that is fully polynomial. We present here an improved version of the model and we comment on the original model in Appendix A.2.
EEF-CYCLE is based on $|\mathcal{R}|$ subgraphs of the compatibility graph. For every $s\in\mathcal{R}$, subgraph $\mathcal{G}^{s}=(\mathcal{R}^{s},\mathcal{A}^{s})$ is the graph induced by vertex set $\mathcal{R}^{s}=\{v\in\mathcal{R}:v\geq s\}$, thus having arc set $\mathcal{A}^{s}=\{(u,v)\in\mathcal{A}_{\mathcal{R}}:u,v\in\mathcal{R}^{s}\}$ [^6] Note that one could also replace each subgraph $\mathcal{G}^{s}$ by the full compatibility graph $\mathcal{G}$, but this would result in an unnecessarily large number of variables.. The subgraphs allow us to efficiently exclude cycles of length more than $K$ by limiting the number of selected arcs per subgraph to at most $K$. Namely, by introducing a binary decision variable $x_{uv}^{s}$ for every arc $(u,v)\in\mathcal{A}^{s}$ in every subgraph $s\in\mathcal{R}$, taking value $1$ if arc $(u,v)$ is selected in subgraph $s$, and value 00 otherwise, we obtain the following model:
[…]
@@
The objective function (31) maximises the total weight and constraints (32) enforce that all RDPs are involved in at most one cycle. Moreover, flow conservation constraints (33) ensure that if any vertex in any subgraph has an incoming flow, then it must also have an outgoing flow, which makes sure that all selected arcs form a set of cycles. Finally, constraints (34) limit the number of selected arcs per subgraph to at most $K$, which guarantees that the maximum cycle length is never exceeded. These constraints also break symmetry by requiring that for every subgraph $s\in\mathcal{R}$, if at least one arc is selected in that subgraph, then at least one such arc must leave vertex $s$ in that subgraph. Note that when $K\geq 2$ (or $K\geq 4$ if no compatible RDPs are considered), it is possible that multiple cycles are selected in a subgraph. However, that does not present a problem, as all selected cycles will still satisfy the cardinality constraint.
Example 1.
(continued) Consider the instance presented in Figure 3 for $K=4$. The subgraphs $\mathcal{G}^{1}$, $\mathcal{G}^{2}$, $\mathcal{G}^{3}$ and $\mathcal{G}^{4}$ required by EEF-CYCLE are presented in Figure 4. Overall, there are $10$ arcs across the subgraphs, leading to $10$ variables. For instance, cycle $\left\langle 2,3,4,2\right\rangle$ is obtained by setting $x_{23}^{2}=x_{34}^{2}=x_{42}^{2}=1$. Note that even though these arcs are present in subgraph $\mathcal{G}^{1}$ too, any solution with $x_{23}^{1}=x_{34}^{1}=x_{42}^{1}=1$ is forbidden due to constraints (34), as no arc leaving vertex $1$ is selected in that subgraph.
1234
1234
1234
1234
@@
When adapting EF-CYCLE to chains, the key idea is to construct again subgraphs of the compatibility graph, namely one subgraph $\mathcal{G}^{s}$ per NDD $s\in\mathcal{N}$. We consider two chain models based on EEF-CYCLE. The first, called EEF-CHAIN-EXP, is a variant of a model proposed by Anderson et al., (2015), whereas the second, called EEF-CHAIN-MTZ, is a new adaptation that includes the timestamp variables as proposed by Mak-Hau, (2017) in the context of the model EF-CHAIN-MTZ. However, as the EEF-based chain models are relatively straightforward adaptations of EEF-CYCLE (using similar ideas as the EF-based chain models), we present EEF-CHAIN-EXP and EEF-CHAIN-MTZ in Appendix A.1.
@@
Regarding model-related improvements, Constantino et al., (2013) proposed a pre-processing algorithm for graph reduction similar to those proposed for the EF-based models, but per subgraph. That is, in EEF-CYCLE, for every subgraph $\mathcal{G}^{s}$ for all $s\in\mathcal{R}$, we can remove all vertices/arcs that cannot appear in any cycle of length at most $K$ having $s$ as the lowest-indexed vertex. Similarly, in EEF-CHAIN-EXP and EEF-CHAIN-MTZ, for every subgraph $\mathcal{G}^{s}$ for all $s\in\mathcal{N}$, we can remove all vertices/arcs that cannot appear in any chain of length at most $L$ initiated by NDD $s$. The details are presented in Appendix A.5. Furthermore, for EEF-CHAIN-EXP and EEF-CHAIN-MTZ, it is possible again to omit the variables associated with the arcs $(v,\tau)\in\mathcal{A}_{\tau}$ going to $\tau$, the details of which are presented in Appendix A.6.
@@
Moreover, as for HCF-CYCLE, the ordering of the vertices plays an important role for EEF-CYCLE. Delorme et al., (2023) experimentally showed that even though sorting the vertices by descending total degree resulted in smaller models, sorting the vertices by ascending degree resulted in a stronger linear relaxation, which outweighed the downside of having a larger model in their tested instances. Conversely, Arslan et al., (2024) constructed graph copies based on a feedback vertex set. Their procedure can be seen as the determination of a descending ordering of the vertices in a dynamic fashion. On the other hand, the ordering of the vertices does not impact EEF-CHAIN-EXP and EEF-CHAIN-MTZ. Related to this, we mention that Zeynivand et al., (2024) take a different approach for EEF-CYCLE in which rather than trying to reduce the size of the $|\mathcal{R}|$ subgraphs, they focus on reducing the number of required subgraphs. One downside of their approach is that it becomes harder to reduce symmetry. For example, one can no longer add the factor $\sum_{v:(s,v)\in\mathcal{A}^{s}}x_{sv}^{s}$ on the RHS of constraints (34).

### v2

@@ 3.5 Edge Formulation
In EF-CYCLE, we introduce a binary decision variable $x_{uv}$ for every arc $(u,v)\in\mathcal{A}_{\mathcal{R}}$, taking value $1$ if arc $(u,v)$ is selected, and value 0 otherwise. Moreover, we define $\mathcal{P}_{K-1}$ as the set of all maximal cycle-feasible paths, where a maximal cycle-feasible path $p$ is a simple path in $\mathcal{G}$ of length K1K-1. We denote the $i$th vertex of path $p$ by $p_{i}$ for $i=1,\ldots,K$. Note that it is only possible to select all the arcs along a maximal cycle-feasible path if the arc $(p_{K},p_{1})$ from the head of the path to the tail of the arc (if it exists) is selected as well. This leads to the following definition of EF-CYCLE:
$(\texttt{EF-CYCLE})\quad$ $\max$ $\sum_{(u,v)\in\mathcal{A}_{\mathcal{R}}}w_{uv}x_{uv}$ (16)
@@
The objective function (16) maximises the total weight and constraints (17) enforce that each RDP is involved in at most one cycle. Moreover, constraints (18) are the flow conservation constraints ensuring that whenever a unit of flow enters a vertex, it must leave the vertex too (i.e., if a recipient receives a kidney, their paired donor should donate a kidney as well), and constraints (19) are the long-cycle elimination constraints, which forbid all cycles with length exceeding KK [^9] Strictly speaking, the term $x_{p_{K},p_{1}}$ should be multiplied by $\mathbbm{1}_{(p_{K},p_{1})\in\mathcal{A}_{\mathcal{R}}}$ to ensure that it is only subtracted if the underlying arc $(p_{K},p_{1})$ exists. For brevity, we omit these coefficients in this model and in the models that follow, in analogous scenarios..
In the following, we discuss two chain models based on EF-CYCLE, which we call EF-CHAIN-EXP and EF-CHAIN-MTZ. The former is strongly inspired by models introduced by Constantino et al., (2013) and Anderson et al., 2015b , whereas the latter is a variant of a model introduced by Mak-Hau, (2017). Mak-Hau, (2017) combined their version of EF-CHAIN-MTZ with both EF-CYCLE and EEF-CYCLE, leading to the “exponential-sized SPLIT formulation” and the “polynomial-sized SPLIT formulation”. The main challenge in adapting EF-CYCLE to a chain model is that not only all chains of length more than $L$ must be excluded, but also that all cycles must be forbidden (following our definition of chain models). To deal with this, EF-CHAIN-EXP contains an exponential number of constraints. On the other hand, EF-CHAIN-MTZ borrows ideas from the Miller-Tucker-Zemlin model for the travelling salesman problem (Miller et al., , 1960), resulting in a model of polynomial size.
In both EF-CHAIN-EXP and EF-CHAIN-MTZ, we introduce a binary decision variable $y_{uv}$ for every arc $(u,v)\in\mathcal{A}$ (including not only $\mathcal{A}_{\mathcal{R}}$, but also $\mathcal{A}_{\mathcal{N}}$ and $\mathcal{A}_{\tau}$), taking value $1$ if arc $(u,v)$ is selected, and value 0 otherwise. For EF-CHAIN-EXP, we define $\mathcal{P}^{\prime}_{L-1}$ as the set of minimal chain-infeasible paths, where a minimal chain-infeasible path $p$ is a simple path in $\mathcal{G}$ of length L1L-1 restricted to the vertices in $\mathcal{R}$. Note that it is not possible to use all the arcs along a minimal chain-infeasible path, as this would result in a chain of length exceeding $L$ (after initialising the chain by an NDD and terminating the chain in $\tau$). Moreover, we define $\mathcal{C}_{\leq L-1}$ as the set of cycles of length at most L1L-1. EF-CHAIN-EXP is then as follows:
$(\texttt{EF-CHAIN-EXP})\quad$ $\max$ $\sum_{(u,v)\in\mathcal{A}}w_{uv}y_{uv}$ (21)
@@
The objective function (21) maximises the total weight, constraints (22) and (23) enforce , respectively, that each NDD and each RDP is involved in at most one chain, and constraints (24) ensure flow conservation. Furthermore, constraints (25) forbid all chains of length more than $L$. Note that these constraints automatically also forbid all cycles of length at least $L$. Therefore, constraints (26) are added to forbid all the other cycles (i.e., those with length at most L1L-1). We mention that Constantino et al., (2013) and Anderson et al., 2015b consider a slightly different setting. In particular, the model proposed by Constantino et al., (2013) does not contain cycle-elimination constraints, whereas the model by Anderson et al., 2015b does not eliminate the long chains.
On the other hand, the model EF-CHAIN-MTZ does not involve sets $\mathcal{P}^{\prime}_{L-1}$ and $\mathcal{C}_{\leq L-1}$, but does contain an additional “timestamp” variable $t_{v}$ for every RDP $v\in\mathcal{R}$, resulting in the following model [^10] The term $(L-3)y_{vu}$ in constraints (29) is omitted when $L\leq 2$.:
$(\texttt{EF-CHAIN-MTZ})\quad$ $\max$ $\sum_{(u,v)\in\mathcal{A}}w_{uv}y_{uv}$ (28)
@@
Constraints (29) and (30) together ensure that the maximum chain length is respected and that all cycles are forbidden. Indeed, for each arc $(u,v)\in\mathcal{A}_{\mathcal{R}}$, the associated constraint from (29) imposes that $t_{v}\geq t_{u}+1$ if the arc is selected (i.e., if $y_{uv}=1$), whatever the value of $y_{vu}$. This means that the $t$-variables can really be interpreted as timestamps. Moreover, the term $(L-3)y_{vu}$ (which is only added when the underlying arc $(v,u)$ exists and $L\geq 3$), is optional, but strengthens the LP relaxation. Indeed, for every $(u,v)\in\mathcal{A}_{\mathcal{R}}$, constraints (29) for $(u,v)$ and $(v,u)$ together imply that $y_{uv}+y_{vu}\leq 1$, which can be used to deduce that $y_{vu}=0$ when $y_{uv}=1$, which in turn implies that $t_{v}=t_{u}+1$. Mak-Hau, (2017) presented an equivalent formulation that includes additional timestamp variables $t_{v}$ for all $v\in\mathcal{N}\cup\{\tau\}$.
@@
Regarding model-related improvements, Mak-Hau, (2017) proposed reducing the model size by removing vertices and arcs that cannot appear in any solution respecting the cardinality constraints. That is, for EF-CYCLE we can remove all vertices/arcs that cannot appear in any cycle of length at most $K$, and in EF-CHAIN we can remove all vertices/arcs that cannot appear in any chain of length at most $L$. For completeness, we present these pre-processing algorithms in Appendix F. Furthermore, for EF-CHAIN-EXP and EF-CHAIN-MTZ, it is also possible to omit the variables $y_{v\tau}$ for arcs $(v,\tau)\in\mathcal{A}_{\tau}$, assuming that all weights $w_{v\tau}$ for arcs $(v,\tau)\in\mathcal{A}_{\tau}$ are nonnegative, which we describe in detail in Appendix G.
@@
Moreover, note that the number of constraints in EF-CYCLE and EF-CHAIN-EXP is exponential in $K$ and $L$, respectively, which can be problematic. This concerns constraints (19) for EF-CYCLE and constraints (25) and (26) for EF-CHAIN-EXP. Nevertheless, Constantino et al., (2013) and Mak-Hau, (2017) showed that enumerating all constraints is a viable approach in certain cases (i.e., when the compatibility graph is sufficiently small and sparse and $K$ and $L$ are sufficiently small too). However, in the experiments by Constantino et al., (2013), there were for example already too many constraints for EF-CYCLE for high-density instances with $|\mathcal{R}|=20$ and $K\geq 5$. Alternatively, one can apply constraint generation, in which the problematic constraints are initially removed, after which one iteratively only adds the constraints that have been violated by previous incumbent (or fractional, depending on the choice of implementation) solutions until an optimal solution is found. This was for example applied by Abraham et al., (2007), Anderson et al., 2015b , Mak-Hau, (2018) and Delorme et al., (2023).
@@
Furthermore, we mention that research has been done on alternatives for some of the problematic constraints mentioned above. In Appendix B we discuss two of the alternatives that were proposed for constraints (19) in EF-CYCLE and one alternative that was proposed for constraints (26) in EF-CHAIN-EXP. However, our preliminary computational experiments indicated that these alternatives either performed similarly to (or worse than) the versions that we presented above. Furthermore, we note that in theory, when EF-CHAIN-EXP is combined with a cycle model, we only need the constraints (26) that forbid cycles of length more than $K$. This would allow us to reduce the number of constraints, at the cost of introducing some symmetry (as some cycles can then be obtained through either the cycle component of the model or the chain component of the model). For simplicity, we do not consider this in our computational experiments.
@@
Last, as mentioned in Section 3.2, it is possible to model cycles and chains concurrently with a single set of variables using ideas similar to those underlying EF-CYCLE and EF-CHAIN-EXP. Such a hybrid model has the potential advantage that its size may be smaller than that of the combined model composed of EF-CYCLE and EF-CHAIN-EXP. However, whereas in a combined model with two separate sets of variables it is relatively easy to forbid all cycles of length more than $K$ without excluding any feasible chain, while also forbidding all chains of length more than $L$ without excluding any feasible cycle, this is a real challenge for a hybrid model with a single set of variables. Indeed, Constantino et al., (2013) present a model that only applies when $L\leq K+1$, as their constraints that eliminate long cycles also forbid feasible chains, otherwise. Alternatively, Anderson et al., 2015b avoided part of the problem by presenting a model for the case where $L=\infty$. They also explained how their model could be extended to the case where $L$ is finite. However, their underlying idea is more similar to the idea behind the chain model EEF-CHAIN-EXP, which we present in Section 3.6. On the other hand, our new model, denoted by EF-HYBRID, fits more closely the paradigm of the models EF-CYCLE and EF-CHAIN-EXP that were presented earlier in this section. However, our experiments showed that EF-HYBRID does not perform well (see Section 4.2.1). Therefore, we present this model in Appendix C.
@@ 3.6 Extended Edge Formulation
The main downsides of the models presented in Sections 3.3-3.5 are the exponential number of variables or constraints. On the other hand, the model EEF-CYCLE, introduced by Constantino et al., (2013), can be seen as an extended formulation based on EF-CYCLE that is fully polynomial. We present here an improved version of the model and we comment on the original model in Appendix B.
EEF-CYCLE is based on $|\mathcal{R}|$ subgraphs of the compatibility graph. For every $s\in\mathcal{R}$, subgraph $\mathcal{G}^{s}=(\mathcal{R}^{s},\mathcal{A}^{s})$ is the graph induced by vertex set $\mathcal{R}^{s}=\{v\in\mathcal{R}:v\geq s\}$, thus having arc set $\mathcal{A}^{s}=\{(u,v)\in\mathcal{A}_{\mathcal{R}}:u,v\in\mathcal{R}^{s}\}$ [^11] Note that one could also replace each subgraph $\mathcal{G}^{s}$ by the full compatibility graph $\mathcal{G}$, but this would result in an unnecessarily large number of variables.. The subgraphs allow us to efficiently exclude cycles of length more than $K$ by limiting the number of selected arcs per subgraph to at most $K$. Namely, by introducing a binary decision variable $x_{uv}^{s}$ for every arc $(u,v)\in\mathcal{A}^{s}$ in every subgraph $s\in\mathcal{R}$, taking value $1$ if arc $(u,v)$ is selected in subgraph $s$, and value 0 otherwise, we obtain the following model:
$(\texttt{EEF-CYCLE})\quad$ $\max$ $\sum_{s\in\mathcal{R}}\sum_{(u,v)\in\mathcal{A}^{s}}w_{uv}x_{uv}^{s}$ (31)
@@
The objective function (31) maximises the total weight and constraints (32) enforce that each RDP is involved in at most one cycle. Moreover, flow conservation constraints (33) ensure that if any vertex in any subgraph has an incoming flow, then it must also have an outgoing flow, which makes sure that all selected arcs form a set of cycles. Finally, constraints (34) limit the number of selected arcs per subgraph to at most $K$, which guarantees that the maximum cycle length is never exceeded. These constraints also break symmetry by requiring that for every subgraph $s\in\mathcal{R}$, if at least one arc is selected in that subgraph, then at least one such arc must leave vertex $s$ in that subgraph. Note that when $K\geq 2$ (or $K\geq 4$ if no compatible RDPs are considered), it is possible that multiple cycles are selected in a subgraph. However, that does not present a problem, as all selected cycles will still satisfy the cardinality constraint.
@@
When adapting EF-CYCLE to chains, the key idea is to construct again subgraphs of the compatibility graph, namely one subgraph $\mathcal{G}^{s}$ per NDD $s\in\mathcal{N}$. We consider two chain models based on EEF-CYCLE. The first, called EEF-CHAIN-EXP, is a variant of a model proposed by Anderson et al., 2015b , whereas the second, called EEF-CHAIN-MTZ, is a new adaptation that includes the timestamp variables as proposed by Mak-Hau, (2017) in the context of the model EF-CHAIN-MTZ. However, as the EEF-based chain models are relatively straightforward adaptations of EEF-CYCLE (using similar ideas as the EF-based chain models), we present EEF-CHAIN-EXP and EEF-CHAIN-MTZ in Appendix A.
@@
Regarding model-related improvements, Constantino et al., (2013) proposed a pre-processing algorithm for graph reduction similar to those proposed for the EF-based models, but per subgraph. That is, in EEF-CYCLE, for each $s\in\mathcal{R}$, we can remove all vertices/arcs from subgraph $\mathcal{G}^{s}$ that cannot appear in any cycle of length at most $K$ having $s$ as the lowest-indexed vertex. Similarly, in EEF-CHAIN-EXP and EEF-CHAIN-MTZ, for each $s\in\mathcal{N}$, we can remove all vertices/arcs from subgraph $\mathcal{G}^{s}$ that cannot appear in any chain of length at most $L$ initiated by NDD $s$. The details are presented in Appendix F. Furthermore, for EEF-CHAIN-EXP and EEF-CHAIN-MTZ, it is possible again to omit the variables associated with the arcs $(v,\tau)\in\mathcal{A}_{\tau}$ going to $\tau$, the details of which are presented in Appendix G.
@@
Moreover, as for HCF-CYCLE, the ordering of the vertices plays an important role for EEF-CYCLE. Delorme et al., (2023) experimentally showed that even though sorting the vertices by descending total degree resulted in smaller models, sorting the vertices by ascending degree resulted in a stronger LP relaxation, which outweighed the downside of having a larger model in their tested instances. Conversely, Arslan et al., (2024) constructed graph copies based on a feedback vertex set. Their procedure can be seen as the determination of a descending ordering of the vertices in a dynamic fashion. On the other hand, the ordering of the vertices does not impact EEF-CHAIN-EXP and EEF-CHAIN-MTZ. Related to this, we mention that Zeynivand et al., (2024) take a different approach for EEF-CYCLE in which rather than trying to reduce the size of the $|\mathcal{R}|$ subgraphs, they focus on reducing the number of required subgraphs. One downside of their approach is that it becomes harder to reduce symmetry. For example, one can no longer add the factor $\sum_{v:(s,v)\in\mathcal{A}^{s}}x_{sv}^{s}$ on the RHS of constraints (34).